### v1

Abstract
In recent years, federated learning (FL) has emerged as a promising approach for machine learning (ML) and data science across distributed edge devices. With the increasing popularity of FL, resource contention between multiple FL jobs training on the same device population is increasing as well. Scheduling edge resources among multiple FL jobs is different from GPU scheduling for cloud ML because of the ephemeral nature and planetary scale of participating devices as well as the overlapping resource requirements of diverse FL jobs. Existing resource managers for FL jobs opt for random assignment of devices to FL jobs for simplicity and scalability, which leads to poor performance.
In this paper, we present Venn, an FL resource manager, that efficiently schedules ephemeral, heterogeneous devices among many FL jobs, with the goal of reducing their average job completion time (JCT). Venn formulates the Intersection Resource Scheduling (IRS) problem to identify complex resource contention among multiple FL jobs. Then, Venn proposes a contention-aware scheduling heuristic to minimize the average scheduling delay. Furthermore, it proposes a resource-aware device-to-job matching heuristic that focuses on optimizing response collection time by mitigating stragglers. Our evaluation shows that, compared to the state-of-the-art FL resource managers, Venn improves the average JCT by up to 1.88×1.$8$.
footnotetext: 1 Computer Science and Engineering, University of Michigan, Michigan, USA 2 Computer Science and Engineering, University of Illinois at Urbana-Champaign, Illinois, USA. Correspondence to: Jiachen Liu <amberljc@umich.edu>.
1 Introduction
Federated learning (FL) enables distributed edge devices to perform collaborative machine learning (ML) without moving raw data into the cloud Bonawitz et al. (2019); Ramage & Mazzocchi (2020). FL has been adopted by many large corporations including Apple, Meta, Google, and LinkedIn to protect user data privacy while improving user experience. For example, Google adopts FL for a wide range of applications such as speech recognition Warden (2018), healthcare study Sadilek et al. (2021), next-word prediction Hard et al. (2019); Xu et al. (2023), emoji prediction Ramaswamy et al. (2019), and query suggestion on keyboard Yang et al. (2018). Each FL training job in practice often requires 1000$\sim$10000 device participants in each training round and takes 4$\sim$8 days to finish Yang et al. (2018). As the number of FL applications continues to grow, efficient edge resource management has become the key to fast and resource-efficient FL.
As opposed to traditional ML training in the cloud Gu et al. (2019); Xiao et al. (2018); Narayanan et al. (2020), resource management in FL introduces several unique characteristics. First and foremost, the availability of FL resources varies rapidly and unpredictably during the lifetime of an FL job, because client devices may become available, join training, or drop off at any time. Meaning, in addition to job-level unpredictability (e.g., when a training job will converge and finish), we must mitigate device-level unpredictability too.
Second, in comparison to cloud ML, FL involves orders-of-magnitude larger resource pools that are also highly heterogeneous in terms of their hardware capacity, software versions, and training data availability. Even when running on the same population, different FL jobs often compete for different subsets of devices, influenced by their specific model characteristics and training objectives. As such, complex multi-resource contentions are common among multiple ongoing FL jobs, where eligible resources for one job may overlap, contain, or be within those of one or more other jobs.
Figure 1: Composition of the completion time of one round of an FL job.
Finally, due to the collaborative and distributed nature of FL, the job completion time (JCT) of an FL job is significantly affected by both scheduling delay and response collection time (Figure 1). During each training round, the former refers to the time needed to acquire all resources required by the FL job. In contrast, the latter refers to the time needed to collect sufficient responses so that the round can be considered complete. A joint optimization of both these factors is necessary to optimize an FL job’s JCT, but it is inherently challenging due to the trade-off between the two objectives. On the one hand, reducing response collection time often involves cherry-picking ideal devices based on additional criteria, which could prolong the scheduling delay of the job. On the other hand, coarse-grained client selection that promptly assigns all incoming eligible devices prevents swift and qualified responses, which eventually takes more time to collect enough satisfactory responses. Importantly, jointly optimizing job scheduling and job-to-device matching is a well-known NP-hard problem Gu et al. (2019), and it is further magnified by the scale of the devices and the FL-specific contention.
In the early days of FL research Kairouz et al. (2019), most solutions focused on optimizing the response collection time of individual jobs, while ignoring the scheduling delay. With the proliferation of FL in production, large corporations such as Apple Paulik et al. (2021), Meta Huba et al. (2022), and Google Bonawitz et al. (2019) are running many FL jobs concurrently. As a result, they have developed their own FL infrastructures to coordinate multiple FL jobs at production scale. However, despite low-level differences, these FL resource managers can all be described simply as random device-to-job matching in various forms. We show that such random assignment can lead to higher scheduling delays and response collection times as more and more FL jobs vie for a shared device population. Note that prior works on resource allocation with job-specific constraints Thinakaran et al. (2017) are not suitable here either, because they were designed for datacenter resources that are not ephemeral, nor so numerous.
In this paper, we present Venn, an FL resource manager that minimizes the average JCT of multiple FL jobs competing over a large number of heterogeneous devices in two steps. First, Venn addresses the intricate resource contention among FL jobs by formulating it as an Intersection Resource Scheduling (IRS) problem (§4.2), where the resources that one job contends for may overlap, contain, or be within those resources of other jobs. Venn then introduces a contention-aware scheduling heuristic that prioritizes small jobs requiring scarce resources to minimize the average scheduling delay. Second, Venn employs a resource-aware device-to-job matching heuristic to reduce response collection time (§4.3). Together, they jointly optimize the average JCT for a diverse set of FL jobs operating under dynamically changing and uncertain resource conditions.
We have implemented and evaluated Venn across various FL workloads derived from real-world scenarios (§5). Compared to state-of-the-art FL resource allocation solutions Bonawitz et al. (2019); Huba et al. (2022); Paulik et al. (2021), Venn improves the average JCT by up to 1.88×1.$8$.
Overall, we make the following contributions in this paper:
  1. 1.
    We introduce Venn, an FL resource manager designed to enable efficient sharing of heterogeneous devices across a large number of FL jobs.
  2. 2.
    To minimize the average JCT of FL jobs, we propose a scheduling and matching joint solution to optimize both the scheduling delay and response collection time.
  3. 3.
    We have implemented and evaluated Venn, along with its scheduling and matching algorithms, demonstrating improvements in the average JCT compared to the state-of-the-art across various real-world FL workloads.
2 Background and Motivation
2.1 Federated Learning
FL has been widely adopted by the industry to train ML models on larger and more diverse on-device datasets without having to copy raw data into the cloud. Example applications include healthcare study, speech recognition, next-word prediction, etc.. FL is unique both in terms of resources it uses and jobs that use it, which pose unique challenges for FL resource managers.
(a) Diurnal device availability.
(b) Device hardware heterogeneity.
Figure 2: FL resources exhibit both high variance in availability and capacity.
(a) Setup.
(b) Random Matching (6.33).
(c) SRSF (5).
(d) Optimal (4.33).
Figure 3: Toy example of resource allocation for three schedules (with tie-breaking based on job ID) across jobs with varying resource demands and device eligibility requirements. Numbers in parentheses are the average JCTs. Devices check in at a constant rate. Eligible devices only for Emoji jobs are marked with blue; all devices are eligible for the Keyboard job. The label of each client indicates its job assignment. Random Matching and SRSF inefficiently allocate scarce Emoji-eligible devices to Job 1, which already has ample Keyboard-eligible resources. Conversely, the optimal schedule smartly allocates these scarce resources to Jobs 2 followed by Job 3, thereby minimizing average request completion time.
FL Resources
FL resources often refer to edge devices with limited capacities, including smartphones, laptops, and Internet of Things (IoT) devices. Compared to resources used by traditional ML cluster (e.g., GPUs), FL resources are unique in terms of availability and heterogeneity.
Dynamic availability. As FL moves ML computations to user devices, the availability of edge devices for sufficiently long durations become crucial. We analyze a real-world client availability trace Lai et al. (2021a), which encompass 180 million trace items of devices behavior over a week. Figure 1(a) shows that the number of available devices (charging and connected to WiFi) changes over time with diurnal pattern. Hence, there is an inherent scheduling delay for an FL job to acquire a sufficient number of device resources.
Device heterogeneity. The large number of devices involved in FL are heterogeneous in multiple dimensions including data availability, software version, and hardware capacity, which leads to varying response time among devices. Figure 1(b) showcases this heterogeneity, focusing on variations in memory and CPU capacities among edge devices, based on data from Ignatov et al. (2019). It also annotates the minimum hardware requirements needed to execute three popular on-device ML models within a reasonable time. As jobs may have different learning objectives, they may compete for different subsets of devices. When more factors like data availability and software version are considered, these subsets can exhibit relationships that are inclusive, overlapping, or nested, leading to complex contention patterns.
FL Jobs
An FL job is executed in multiple rounds that run sequentially for synchronous FL training. Each round begins by issuing a resource request including its demand size and device specifications. Once the request is sent, that particular round of the job enters a waiting phase where it continuously dispatches tasks to participants and collects their responses until a sufficient number of qualified responses are gathered. This process repeats for each round until all the rounds and therefore the job completes (e.g., it reaches a target accuracy or completes a pre-defined number of rounds).
In this paper, we focus only on synchronous FL jobs.
2.2 FL Resource Managers
To orchestrate large FL training jobs, several FL resource managers have been proposed in the industry with three primary designs that all support synchronous FL jobs.
  1. 1.
    Apple’s FL resource manager Paulik et al. (2021) is driven by clients, where each client independently samples from a list of FL jobs they are able to execute.
  2. 2.
    Meta’s FL resource manager Huba et al. (2022) is centralized, where it randomly matches each client with one eligible FL job.
  3. 3.
    Google’s FL resource manager Bonawitz et al. (2019) is driven by jobs, where each job independently samples from available clients.
Despite the seeming variety in their designs, existing resource managers all boil down to random device-to-job matching in different forms. Unfortunately, in the absence of extreme surplus of qualified devices compared to the demand, random matching fail to address the growing resource contention and yield longer average scheduling delay.
Limitations of the state-of-the-art.
Consider the toy example in Figure 3 that compares three centralized scheduling solutions: random matching as specified above, Shortest-Remaining-Service-First (SRSF) proposed for cloud ML scheduling Gu et al. (2019), and the optimal solution, assuming all jobs arrive simultaneously. Both random matching and SRSF fail to account for resource scarcity and contention, thereby allocating limited resources (i.e., Emoji clients) to jobs (i.e., the Keyboard job) that already have an abundance of resources, resulting in an average JCT of 6.33 and 5 time units. Note that the contention patterns in real-world scenarios are often more complex and larger in scale than the one presented in this example.
2.3 Opportunities for Improvements
Figure 4: Impact of resource contention.
Figure 4: Impact of resource contention.
Figure 5: JCT breakdown in a single round.
Impact of resource contention
Most of existing FL solutions that focus on single-job optimizations, such as optimizing for client selection Lai et al. (2021b); Abdelmoniem et al. (2023); Balakrishnan et al. (2022), often rely on the assumption that FL jobs have access to sufficient number of online devices. However, in practice, there could be multiple FL jobs running at the same time and competing for the same set of devices, leading to resource contention Bonawitz et al. (2019). This can significantly impact the performance of FL jobs, such as accuracy and wall clock time.
We analyze the impact of resource contention on FL jobs’ performance in Figure 5. In this experiment, the resource pool is evenly partitioned and managed by each job, who aims to train a ResNet-18 model for the FEMNIST dataset Cohen et al. (2017). We vary the number of concurrently running jobs to observe its impact. As more jobs share the same device pool, the available device choices for each job become increasingly constrained, leading to a noticeable degradation in the round-to-accuracy performance. Hence, evenly partitioned resource pool for each FL job is no better than a shared device pool in terms of participant diversity.
Breakdown of request completion time
While most existing FL efforts, such as quantization Reisizadeh et al. (2020) and client selection He et al. (2023), predominantly focus on optimizing response collection time—–i.e., the time needed to collect a sufficient number of responses—–they often overlook a critical component: scheduling delay, as depicted in Figure 1.
To offer a holistic view of JCT, we dissect its components for a single round request, under varying degrees of resource contention, as shown in Figure 5. Utilizing the same experimental setup as in Figure 5, we quantify both the average scheduling delay and response collection time during one round of training and testing, involving 100 participants. The scheduling delay is specifically captured under random device-to-job matching, and the shaded regions cover the duration for each job. As our results in Figure 5 indicate, scheduling delay can significantly impact overall JCT, especially when resource supply falls short of demand.
3 Venn Overview
Venn serves as a standalone FL resource manager that operates at a layer above all FL jobs, and it is responsible for allocating each checked-in resource to individual jobs. Figure 6 provides an overview of Venn’s workflow, along with its role in an FL job’s lifecycle.
Figure 6: Venn System Overview.
In each execution round, an FL job submits resource requests to Venn, specifying its device requirements and associated resource demands (0⃝). Devices continuously check in with Venn as they become available over time (1⃝). Based on the real-time resource demand and supply, Venn generates a resource allocation plan to assign one FL job to each checked-in device (2⃝), until the job’s demand is satisfied. Upon receiving the task assignment from Venn, each device adheres to the allocation plan and participates in the corresponding job (3⃝). The device then retrieves the computation plan from the job and performs on-device computation (Lv et al., 2022; CoreML, ) (3⃝ and 4⃝). Finally, the device may either report the training result to the corresponding job upon completion, or drop off mid-process due to availability dynamics (5⃝).
Note that Steps 3⃝ through 5⃝ adhere to conventional FL protocols between individual FL jobs and devices. Venn’s primary role is in optimizing the job-to-device assigning phase, denoted by 2⃝.
Venn delegates responsibilities such as device selection, device fault tolerance, and privacy protection to individual FL jobs. Device failures are both inevitable and difficult to predict in FL. Rather than imposing a one-size-fits-all solution, Venn empowers FL jobs to take the reins on fault tolerance based on their specific workloads and objectives. Therefore, Venn offloads handling device fault tolerance to FL jobs, who can better detect and react to device failures (e.g., deciding the amount of overcommit Bonawitz et al. (2019)). Similarly, Venn offers FL jobs the freedom to design their own device selectors Lai et al. (2021b), where they can incorporate customized resource criteria into their requests. Venn also does not interfere with job-specific privacy solutions such as secure aggregation Bonawitz et al. (2016); Huba et al. (2022) or differential privacy Geyer et al. (2017); Xu et al. (2023).
4 Resource Scheduling in Venn
In this section, we first introduce the problem statement (§4.1). We then develop our scheduling algorithm in two steps: establishing the job scheduling order to minimize the scheduling delay (§4.2) and determining the device-to-job matching to co-optimize the response collection time (§4.3). Then, we discuss additional enhancements for real-world deployments (§4.4).
4.1 Problem Statement
Given a collection of FL jobs—along with their device requirements and resource demands—and a set of heterogeneous devices that arrive and depart over time, Venn should efficiently assign devices to FL jobs in order to reduce the average job completion time (JCT). The scheduling problem can be mathematically modeled as a multi-commodity flow (MCF) problem with integer constraints, where each FL job is modeled as a distinct commodity and each device serves as an intermediate vertex between the source and sink of its corresponding eligible FL job. Then the goal of this integer MCF problem is to minimize the average JCT of jobs, which is known to be NP-hard (Even et al., 1975). Even for its linear approximation, the time complexity is exacerbated by the planetary scale of devices involved and diverse device requirements from jobs, making existing solutions computationally infeasible. We now formally define such resource scheduling problem.
Problem Definition.
Assume we have $m$ jobs $\mathbb{J}=\{J_{1},J_{2},\allowbreak...,J_{m}\}$ with their resource demands $\mathbb{D}=\{D_{1},D_{2},\allowbreak...,D_{m}\}$. Let $\mathbb{S}=S_{1}\cup S_{2}\cup...\cup S_{n}$ be the available device set, where $S_{k}$ is the eligible device subset in $\mathbb{S}$ that satisfies the device specification of job $J_{i}$, i.e., $f(J_{i})=S_{k}$. The goal is to match each checked-in device $s\in S_{k}$ with one job $J_{i}$, where $f(J_{i})=S_{k}$, $\forall s\in\mathbb{S}$, in order to minimize average JCT, which consists of scheduling delay and response collection time.
Tradeoff between scheduling delay and response collection.
Jointly optimizing scheduling delay and response collection time is non-trivial. Intuitively, to reduce the average scheduling delay, it is desirable to promptly assign each device upon check-in to one eligible job. On the other hand, since the response collection time is usually determined by the final responding device among the target number of participants, it can be minimized by mitigating stragglers and cherry picking ideal devices to jobs. However, such an approach takes a longer time to acquire a certain number of qualified devices, leading to longer scheduling delay.
At its core, Venn aims to find a sweet spot in the trade-off in order to optimize average JCT. Specifically, we decouple the FL resource allocation problem as following two questions:
  1. 1.
    How to decide the FL job scheduling order in order to minimize the average scheduling delay? (§4.2)
  2. 2.
    How to match devices with FL jobs in order to minimize average response collection time while also reducing average JCT? (§4.3)
Algorithm 1 Intersection Resource Scheduling
1:function Venn-Sched( Job Groups $\mathbb{G}$, Devices $\mathbb{S}$ )
$\triangleright$    Sort within job group (§4.2.1).
2:     for $G_{j}$ in $G$ do
3:         Sort $J_{i}$ by $D_{i}$ in ascending order, $\forall J_{i}\in G_{j}$
$\triangleright$    Generate initial allocation $S_{j}^{\prime}$ for each group $G_{j}$.
4:     $S=\cup_{j=1}^{n}S_{j}$
5:     Sort $G_{j}$ by $|S_{j}|$ in ascending order, $\forall G_{j}\in\mathbb{G}$
6:     for $G_{j}$ in $G$ do
7:         $S_{j}^{\prime}=S\cap S_{j},S=S\setminus S_{j}^{\prime}$
$\triangleright$    Allocate resource $S_{j}^{\prime}$ for each group $G_{j}$.
8:     Sort $G_{j}$ by $|S_{j}|$ in descending order, $\forall G_{j}\in\mathbb{G}$
9:     for $G_{j}$ in $G$ do
10:         if  $|S_{j}^{\prime}|>0$
11:              for $G_{k}\in\mathbb{G}:|S_{k}|<|S_{j}|,S_{k}\cap S_{j}\neq\emptyset$ do
12:                  $m_{j}^{\prime},m_{k}^{\prime}$= get-queue-len()
13:                  if  $\frac{m_{j}^{\prime}}{|S_{j}^{\prime}|}>\frac{m_{k}^{\prime}}{|S_{k}|}$
14:                       $S_{j}^{\prime}=S_{j}^{\prime}\cup(S_{j}\cap S_{k})$
15:                       $S_{k}^{\prime}=S_{k}^{\prime}-S_{j}^{\prime}$
16:                  else
17:                       break
18:     return $\{G_{j}[0],S_{j}^{\prime}\},\forall j\in[1,n]$
4.2 Intersection Resource Scheduling (IRS)
We now tackle the first question, minimizing the average scheduling delay. Directly matching jobs to devices can be computationally expensive, especially when dealing with the immense scale of devices and jobs. The challenges posed by this problem are not solely due to its scale either; they are also compounded by the diverse resource requirements of FL jobs. These varying requirements introduce intricate resource contention patterns, further complicating the scheduling process.
To this end, Venn introduces the Intersection Resource Scheduling (IRS) problem to account for this resource contention. Basically, each FL job $J_{i}\in\mathbb{J}$ may compete for a subset of devices $S_{k}\in\mathbb{S}$, denoted as $f(J_{i})=S_{k}$, where these resource subsets can exhibit relationships that are inclusive, overlapping, or nested. We created an integer linear programming (ILP) formulation to optimally allocate resources to minimize scheduling delay (Appendix A) and propose a heuristic to tackle the problem.
To tackle the scale of devices and jobs, Venn aims to determine a job scheduling order, where each checked-in device is assigned to the first eligible job in the order, rather than scattering resources across multiple jobs. Such a fixed job order can minimize the scheduling delay while reducing computational complexity.
With the objective of determining a job scheduling order, Venn first groups jobs $\mathbb{J}$ into Resource-Homogeneous Job Groups $\mathbb{G}=\{G_{1},G_{2},...,G_{n}\}$ by their resource requirements, where each job group $G_{j}=\{J_{i}|f(J_{i})=S_{j},\forall J_{i}\in\mathbb{J}\}_{i=1}^{m_{j}}$ contains all jobs with the same resource requirement. Venn addresses the problem using a two-step approach, with each step occurring at a different level of scheduling granularity, as outlined in Algorithm 1: (i) Determining the job order within a job group to optimize local resource scheduling (§4.2.1). (ii) Then deciding how to merge the job order across job groups to ensure global scheduling efficiency (§4.2.2). Venn invokes Algorithm 1 on job’s request arrival and completion. By breaking down the overall problem into two steps, we further reduce the problem’s complexity without affecting the scheduling efficiency. We provide theoretical insights to show the effectiveness of this problem decomposition in Appendix B.
4.2.1 Intra Job Group Scheduling
Within each job group that shares the same device specifications, Venn prioritizes jobs based on their remaining resource demand (Algorithm 1 line 3). This ordering strategy aims to minimize the intra-group scheduling delay. Prioritizing jobs with smaller remaining resource demands has been shown to be effective in similar scheduling problems Garey et al. (1976). We choose this locally optimal scheduling strategy with the observation that it aligns well with the goal of achieving a globally optimal scheduling order. By default, the remaining resource demand refers to the needs of a single request within one round. However, it can also encompass the total remaining demand for all upcoming rounds, provided such data is available.
4.2.2 Inter Job Group Scheduling
Addressing the scheduling problem across multiple job groups introduces an additional layer of complexity due to the intricate patterns of resource contention. Traditional scheduling algorithms, such as Random Matching and Shortest Remaining Service First (SRSF), are not designed to account for resource contention across job groups, leading to poor average JCT.
An effective scheduler should recognize and adapt to the resource contention patterns among jobs. At a high-level, it should prioritize jobs with scarce resources, i.e., those with stringent device requirements and limited eligible resource options, to avoid being blocked by jobs with more abundant resources. Additionally, when a particular resource type is in high demand, the scheduler should judiciously allocate intersected resources to the job group with a longer queue, as this group contributes more significantly to the average scheduling delay.
To achieve this, Venn allocates the intersected resources across different job groups based on two factors related to average scheduling delay:
  1. 1.
    Amount of eligible resources allocated: the job group with smaller amount of eligible resources may have longer scheduling delay under the same condition.
  2. 2.
    Queue length: the job group with a longer queue length contributes more to the average scheduling delay as more jobs are waiting for the same type of resources.
Algorithm 1 outlines the steps Venn takes to allocate the current resource across job groups. First, Venn initializes resource allocation among job groups by starting to allocate resources to job group with most scarce resources (line 5). This results in an initial allocation plan with no resource sharing across job groups (lines 47), setting the stage for subsequent cross-group allocations.
To determine how to allocate intersected resource across job groups, Venn greedily evaluates whether a job group with more abundant resources should acquire intersected resources from groups with scarcer resources with the objective of minimizing the average scheduling delay. This evaluation starts with the job group possessing the most abundant resources (line 8). If the resources allocated to this group remain unclaimed by other groups (line 10), Venn will decide how much resources from subsequent job groups should be allocated to it. Specifically, Venn prioritizes job groups with longer queue lengths and fewer allocated resources, guided by a ratio that balances the number of affected jobs against the amount of allocated resources (line 13). If this ratio $\frac{m_{j}^{\prime}}{|S_{j}^{\prime}|}$ is larger than the one for the target resource group $\frac{m_{k}^{\prime}}{|S_{k}^{\prime}|}$, Venn reallocates resources accordingly (lines 1415). Otherwise, the algorithm ceases to allocate additional resources to the current job group from remaining groups (line 17). The reason is that if this job group needs more resources, it should first take the resources from job groups with relatively abundant resources.
Function get-queue-len() (line 12) would return the number of jobs $m^{\prime}$ whose JCT would be delayed by potential group prioritization. For example, the affected queue length $m^{\prime}$ may contain jobs from other job groups that have been deprioritized previously. An easier way to approximate $m^{\prime}$ is to use the group queue length. If intersected resources have been decided to be allocated to job group $G_{j}$ over $G_{k}$, Venn accumulates and updates their current resource allocation $S_{j}^{\prime},S_{k}^{\prime}$ and queuing length $m_{j}^{\prime}$ (line 1415).
The time complexity of Algorithm 1 is $\max(O(m\log m),O(n^{2}))$, where $m$ is the number of ongoing jobs and $n$ is the number of job groups. To formalize the effectiveness of Venn’s approach, we illustrate the theoretical insight behind the scheduling algorithm in Appendix C.
4.3 Device Matching
Now we focus on minimizing the response collection time, a significant contributor to the overall JCT, particularly when resource contention is low. Existing cluster-level device-to-job matching solutions, either stick to a certain job order such as FIFO and SRSF, or match devices without a strategic algorithm such as random match, where none of them optimizes the job response collection time.
Response collection time is usually determined by the last successfully responding devices. Hence, it can be reduced by allocating devices with similar higher capacity to the FL job. Meanwhile, these high-end devices have lower probability to fail due to their quick task execution.
Figure 7: Visualize tier-based device-to-job matching condition.
However, as mentioned in Section 4.1, there is trade-off between scheduling delay and response collection time. Intuitively, with limited device influx, priority should be given to minimizing the scheduling delay, which dominates the average JCT. On the other hand, with sufficient device influx to fulfill a job request within a short period, we should consider minimizing the response collection time while obeying the scheduling order given by Section 4.2.
Algorithm 2 Device Matching
1:$S=\{S^{1},S^{2},...,S^{V}\}$triangleright Evenly partition devices
2:$g_{v}=\frac{t^{v}}{t^{0}},\forall v\in[1,V]$triangleright Response time speed up
3:Input: $J_{i}$, $S_{j}^{\prime}\in{Venn}-SCHED(\mathbb{G},\mathbb{S})$
4:function Venn-Match( Job $J_{i}$, Resource $S_{j}^{\prime}$ )
5:     $c_{i}=\frac{t_{response}}{t_{schedule}}$
$\triangleright$     Assign tiers in a rotating manner.
6:     u = randint(0, V)
$\triangleright$     Decide whether to trigger tier-based matching
7:     if $V+g_{u}*c_{i}<c_{i}+1$
$\triangleright$         Assign tier $u$ devices to $J_{i}$
8:         Update $S_{j}^{\prime}=S_{j}^{\prime}\cap S^{u}$
9:     return $\{J_{i},S_{j}^{\prime}\}$
To this end, we propose a resource-aware tier-based device-to-job matching solution to reduce the response collection time for each job as illustrated in Algorithm 2.
The matching algorithm is activated only for jobs that are currently served, as scheduled by Algorithm 1. For each such job, Venn partitions the eligible devices into $V$ tiers based on their hardware capabilities, where $V$ denotes the granularity of this partitioning. If a job has been served before, Venn adaptively sets the tier partition thresholds based on the hardware capacity distribution of the devices that participated in earlier rounds. Otherwise, Venn forgoes tier-based matching and profiles the devices allocated to the job’s current request to inform future device tier partitioning.
For each served job request, Venn randomly selects a device tier, denoted as $S^{u}$, to the job (Algorithm 2 line 6). This randomized tier assignment aims to expose each FL job to a diverse set of devices, rather than confining them to high-end devices. Given that the response collection time is determined by the slowest responding participant, tier-based assignment does not adversely affect this metric.
As illustrated in Figure 7, while tier-based matching may increase the scheduling delay by a factor of $V\geq 1$, it can concurrently reduce the response collection time by a factor of $g\leq 1$. The algorithm proceeds to perform such tier-based device-to-job matching for the job $J_{i}$ only if its JCT can be reduced, i.e., $1+c_{i}>V+c_{i}g_{u}$ (line 7). If the condition holds, Venn allocates device tier $S_{u}$ to the job, effectively updating the set to $S_{j}^{\prime}\cap S^{u}$. Meanwhile the leftover device tiers would be allocated to subsequent jobs in the job group, maximizing the utilization of available resources.
To determine the response time speed-up factor $g$ for tier-based matching, we note that the device response time distribution adheres to a log-normal distribution (Wang et al., 2023). We use the 95th percentile as the statistical tail latency to account for the overall response collection time, thereby excluding failures and stragglers. Venn profiles and estimates the response collection time for each device tier $v\in[1,..,V]$ and subsequently computes the speed-up factor $g_{v}=\frac{t^{v}}{t^{0}}$ relative to a non-tiered scenario (line 2).
4.4 Enhancements
Dynamic Resource Supply
As shown in Figure 1(a), the total available FL resources change significantly over time. To address this, Venn initially records each device eligibility through a time-series database. This database is then queried for resource eligibility distribution from a past window, which is sourced from public device check-in information. However, relying solely on momentary eligible resource rates for input into the scheduling algorithm is inaccurate. This is primarily due to the varying resource arrival patterns, as demonstrated in Figure 1(a). Given that FL jobs often span multiple days and resource availability typically follows a diurnal pattern, a more accurate approach is to use the average eligible resource rate over a 24-hour period as a representative metric for each job’s eligible resources. As a result, the scheduler can become both farsighted and robust, effectively accommodating the dynamic nature of resource availability.
Starvation Prevention
Our heuristic can lead to larger FL jobs being starved due to the preference given to smaller jobs. This is not acceptable especially when the jobs are initiated by different FL developers who require performance guarantees. Venn grants fairness to jobs to avoid such starvation. Specifically, our goal is to guarantee that the scheduling latency of a job $J_{i}$ is no worse than fair sharing, which is defined as $T_{i}=M*sd_{i}$, where $M$ is the number of simultaneous FL jobs and $sd_{i}$ represents the JCT without contention. Then, we adjust each job demand to be $d_{i}^{\prime}=d_{i}\times(\frac{t_{i}}{T_{i}})^{\epsilon}$ to ensure fairness within a job group, and adjust each group queue length $q_{j}^{\prime}=q_{j}\times(\frac{\sum_{J_{i}\in G_{j}}T_{i}}{\sum_{J_{i}\in G_{j}}t_{i}})^{\epsilon}$ to ensure fairness across job groups. $t_{i}$ is the time usage of job $J_{i}$ at the moment and $\epsilon\in[0,\infty)$ is a fairness control knob. When $\epsilon=0$, the algorithm is identical to the one in Section 4.2. As $\epsilon\rightarrow\infty$, the fairness multiplier dominates the scheduling, resulting in maximum fairness. We show that Venn improves JCT over its counterparts with our starvation design (§5.5).
5 Evaluation
In this section, we evaluate the effectiveness of Venn through event-driven simulation and testbed experiments. Our key takeaways are:
  • Venn speeds up the average JCT by up to 1.88X without affecting the model accuracy, compared to the state-of-the-art across various real-world FL workloads (5.2).
  • Venn outperforms its counterparts through intelligent job scheduling and device-to-job matching using different design components (5.3).
  • Venn ’s benefits are robust under a wide range of FL workloads and environment setups (5.5).
5.1 Experiment Setup
(a) Device eligibility trace.
(b) FL job demand trace.
Figure 8: Device and job trace used in experiments. (a) Devices are stratified into four regions to explore different overlap patterns. (b) The diverse workloads in experiments are derived from the job demand trace based on demand characteristics.
Testbed
To rigorously evaluate Venn, we employ a two-pronged approach. First, we have developed a high-fidelity simulator that replays client and job traces, effectively emulating the dynamics of the scheduling environment for large-scale evaluations. Second, we deploy real FL systems to execute actual FL jobs at a smaller scale of devices.
FL Resources:
To faithfully emulate heterogeneous device runtimes, network throughput, and availability, we use device traces from FedScale Lai et al. (2021a) and AI Benchmark Ignatov et al. (2019), as depicted in Figure 2 and Figure 7(a). Each unique device trace is limited to one FL job per day for realism.
FL Jobs:
Our focus is on synchronous FL jobs Yang et al. (2018), where each successful training round requires a minimum of 80% target participants to report back within a deadline, which is set to be 5min - 15min depending on the round demand. To assess the generalizability of Venn, we curate a diverse set of FL jobs drawn from diverse applications Xu et al. (2023); Yang et al. (2018); Hard et al. (2019); Ramaswamy et al. (2019); Warden (2018); Sadilek et al. (2021), whose resource demand is depicted in Figure 7(b). In the real FL experiment, each job aims to train a ResNet-18 He et al. (2016) and MobileNet-V2 Sandler et al. (2018) on FEMNIST dataset.
Workloads:
Our evaluation includes five workload scenarios that sample differently from the job trace in Figure 7(b) to rigorously evaluate Venn’s performance. Even: Sampled from all jobs, which is the default trace. Small: Uniformly sampled only from jobs with below-average total demand. Large: Uniformly sampled only from jobs with above-average total demand. Low: Uniformly sampled only from jobs with below-average demand per round. High: Uniformly sampled only from jobs with above-average demand per round.
Devices are stratified into four specifications based on their CPU and memory capacities (Figure 7(a)) to create various resource contention patterns where the eligible resource set for one job may overlap, contain, or be within the eligible resource set of other jobs. By default, each job is randomly mapped to one device specification among General resources, Compute-Rich resources, Memory-Rich resources, High-Performance resources. Jobs arrive via a Poisson process with a 30-min average inter-arrival. Default simulation and real-world workloads contain 50 and 20 jobs, respectively.
Baselines
We compare Venn with FIFO, SRSF, and an optimized random matching. Both Venn and SRSF are agnostic to the total FL job rounds. Random matching algorithm is supposed to match one eligible job to each device at random. However, to reduce its round abortion rate under contention, we optimize it to schedule job requests in a randomized order, thereby setting a more challenging baseline. Note that we only run Venn with the starvation prevention strategy in Section 5.5.
Metrics:
Our primary performance metrics include the average job completion time (JCT). Note that while Venn does not explicitly optimize for FL job accuracy, it does not adversely affect it either.
  FIFO SRSF Venn
Even 1.38$\times$ 1.69$\times$ 1.87$\times$
Small 1.48$\times$ 1.68$\times$ 1.78$\times$
Large 1.64$\times$ 1.57$\times$ 1.72$\times$
Low 1.55$\times$ 1.66$\times$ 1.88$\times$
High 1.42$\times$ 1.41$\times$ 1.63$\times$
Table 1: Summary of improvements on average JCT over random matching on different FL workloads.
5.2 End-to-End Performance
Venn achieves better average JCT Improvement. We assess the performance of different scheduling algorithms by evaluating its performance over different workloads. We report the average JCT speed-up for each scheduling algorithms compared to the random scheduling in Table 1. We observe that our scheduling algorithm consistently provides stable improvements in the average JCT across various workloads, which underscores the robustness of Venn.
Venn achieves faster convergence without affecting accuracy. We report the final model test accuracy of FL jobs under different schedules with the help of our FL system at a smaller experiment scale. As shown in Figure 10, we observe that Venn does not affect the final model test accuracy but speeds up the overall convergence process.
Venn has negligible overhead. We emulated a large number of FL jobs and groups to evaluate the scheduler’s scalability. Our results in Figure 10 demonstrate that the latency incurred by one-time triggering for scheduling and matching remains low, even with a substantial increase in job and group numbers. This can be attributed to its time complexity $\max(O(m\log m),O(n^{2}))$, where $m$, $n$ are the numbers of ongoing jobs and job groups, respectively.
Figure 9: Venn does not affect the average test accuracy.
Figure 9: Venn does not affect the average test accuracy.
Figure 10: Venn introduces negligible overhead at scale.
5.3 Performance Breakdown
(a) Low workload.
(b) High workload.
Figure 11: Average JCT improvement breakdown.
  25th 50th 75th
Even 11.5$\times$ 7.2$\times$ 5.6$\times$
Small 6.8$\times$ 5.2$\times$ 4.3$\times$
Large 3.7$\times$ 2.9$\times$ 2.7$\times$
Low 11.6$\times$ 7.5$\times$ 4.7$\times$
High 5.1$\times$ 3.3$\times$ 3.1$\times$
Table 2: Breakdown of average JCT Improvement across jobs with lowest 25%, 50%, and 75% of total demands. Venn benefits more on smaller jobs.
  General. Compute. Memory. High-perf.
Even 1.5$\times$ 7.2$\times$ 5.3$\times$ 3.9$\times$
Small 0.9$\times$ 6.0$\times$ 2.8$\times$ 2.6$\times$
Large 0.9$\times$ 3.7$\times$ 1.8$\times$ 2.6$\times$
Low 0.8$\times$ 3.4$\times$ 2.1$\times$ 8.7$\times$
High 0.8$\times$ 2.2$\times$ 2.2$\times$ 5.6$\times$
Table 3: Breakdown of average JCT Improvement across jobs that ask for different resource requirements: General resources, Compute-rich resources, Memory-rich resources and High-performance resources. Venn benefits more on jobs that ask for scarcer resources.
We present a performance breakdown of Venn, which consists of two parts: a job scheduling algorithm that determines the job order to minimize the scheduling delay, and a device-to-job matching algorithm to reduce the response collection time. We evaluate the performance of Venn with only the scheduling algorithm (Venn w/o matching), Venn with only matching algorithm and FIFO (Venn w/o scheduling), and Venn with both algorithms (Venn). We show the improvement of the average JCT over the default random scheduling for each component. As shown in Figure 11, the tier-based device-to-job matching algorithm primarily benefits low workload where the resource contention is small, which is aligned with our original design intention. The reason is that when resource supply is sufficient, the response collection time would dominate the JCT, which can be optimized by our matching algorithm.
To analyze the impact of Venn on different types of jobs, we break down jobs by their total demands and device requirements (Figure 7(a)), and then analyze the average JCT improvement for each type. Table 2 and Table 3 quantifies how Venn improves average JCT across varying total demands (25th, 50th, 75th percentiles) and eligibility types (General, Compute-rich, Memory rich, High-performance ). Notably, jobs that have smaller total demands and ask for scarce resources benefit the most from Venn.
5.4 Case Study on Biased Workload
This section delves into an in-depth analysis of Venn’s adaptability and performance across four distinct workloads, each characterized by a specific bias in job resource requirements. These workloads include General, Compute-Heavy, Memory-Heavy, and Resource-Heavy categories. For example, the Compute-Heavy workload is structured such that half of its jobs are predominantly geared towards compute-intensive resources, with the rest evenly distributed across the other three resource types. This setup introduces varied queue lengths in different job groups, providing a robust testbed for evaluating Venn’s capability in effectively managing these variations.
The design of these workloads aims to scrutinize Venn’s proficiency in navigating diverse resource requirement distributions, while maintaining uniformity in job demands as illustrated in Figure 7(b). The outcomes of these experiments are systematically presented in Table 4, offering insights into the algorithm’s performance under each workload scenario.
  FIFO SRSF Venn
General 1.46$\times$ 1.78$\times$ 1.94$\times$
Compute-heavy 1.73$\times$ 2.08$\times$ 2.23$\times$
Memory-heavy 1.68$\times$ 2.05$\times$ 2.27$\times$
Resource-heavy 1.65$\times$ 1.90$\times$ 2.01$\times$
Table 4: Average JCT improvement on four biased workloads.
5.5 Ablation Study
Figure 12: Venn outperforms FIFO and SRSF across different numbers of jobs.
Figure 12: Venn outperforms FIFO and SRSF across different numbers of jobs.
Figure 13: Venn’s improvement across different numbers of tiers.
Impact of number of jobs:
We evaluate Venn under various scenarios with different numbers of FL jobs arriving over time. As the number of jobs increases, resource contention becomes more pronounced, highlighting the importance of efficient scheduling under such conditions. We present the average JCT speed-up with different numbers of jobs in even workload to demonstrate the effectiveness of our algorithm. As shown in Figure 13, Venn consistently provides improvement across various numbers of jobs, with its benefits becoming more pronounced as the number of jobs increases.
Impact of number of tiers:
We evaluate the matching algorithm’s performance across varying numbers of client tiers, ranging from a single tier to multiple. Figure 13 shows that increased tier granularity enhances device-to-job matching and improves performance. However, the gains plateau beyond a certain point, as finer tiers increase scheduling delay without yielding further reductions in response collection time. Thus, optimizing the number of tiers is crucial for balancing scheduling efficiency and JCT improvement.
(a) Venn’s improvement over different $\epsilon$.
(b) Ratio of jobs meet the fair-share JCT.
Figure 14: Fairness knob.
Impact of fairness knob:
We incorporated a fairness knob ($\epsilon$) to strike a balance between performance and fairness. We conducted an evaluation using various values of $\epsilon$, plotting the average JCT speed-up against these values in Figure 13(a). The results demonstrate that, as $\epsilon$ increases, the JCT speed-up tends to decrease. As shown in Figure 13(b), the percentage of jobs that meet the fair-share JCT increases with the $\epsilon$, where ϵ=italic-$\epsilon=$2 gives 69% of jobs receive their fair-share JCT. This observation highlights the trade-off between performance and fairness within our FL resource scheduling algorithm, which can be fine-tuned by adjusting the value of $\epsilon$.
6 Related Works
Cluster Resource Manager. There are many cluster resource managers that schedule heterogeneous resources with constraints Thinakaran et al. (2017); Ghodsi et al. (2013); Narayanan et al. (2020). However, they are mostly designed for data center and fail to capture the level of availability, heterogeneity and scale of FL resources, nor do they consider both scheduling delay and response collection time in the JCT. Existing ML cluster resource managers mainly focus on managing the stable data-center resources like GPU and CPU (Xiao et al., 2018; Mohan et al., 2022) in order to improve JCT, utilization and fairness (Peng et al., 2018; Chaudhary et al., 2020). Some research delves into GPU-specific optimizations Gu et al. (2019); Yu et al. (2021); Hwang et al. (2021), while others co-design resource managers with the specific characteristics of ML workloads Qiao et al. (2021); You et al. (2023). However, these ML cluster resource managers are not directly applicable to the more uncertain and heterogeneous FL resources, nor to FL job characteristics. Moreover, we want to note that Venn is complementary to these in-cluster schedulers. For example, when the number of rounds of FL jobs is not available, we can leverage Least-Attained Service Gu et al. (2019) to performed Venn scheduling within each Tiresias-determined job queue.
FL Client Selector. Several recent works have studied client selection at the single FL job level. In addition to enforce device requirements including software version, hardware capacity and data quality, they further cherry-pick clients based on their state, system and statistical utility (Lai et al., 2021b; Abdelmoniem et al., 2023; Balakrishnan et al., 2022; Jee Cho et al., 2022; Chai et al., 2020; He et al., 2023) to speed up the training. However, they solely focus on response collection time (Liu et al., 2022a; b) and overlook the time required to acquire adequate resources. Additionally, optimizing individual FL job performance is insufficient as the deployment scale of FL applications continues to grow.
FL Resource Manager. Large companies including Apple, Meta and Google have proposed their FL infrastructures; however, FL resource management is not their primary focus. These resource managers simply adopt random device-to-job matching in various forms, resulting in suboptimal scheduling delays and response collection times.
7 Conclusion
In this paper, we introduce our FL resource manager, Venn, to enable efficient sharing of a large amount of heterogeneous device resources among multiple FL jobs with diverse requirements. Venn incorporates a contention-aware job scheduling algorithm and a resource-aware device-to-job matching algorithm to minimize the average JCT for FL jobs. Our evaluation over various real-world FL workloads shows that Venn achieves up to 1.88X improvement on average JCT.
Appendix A ILP Formulation of IRS
We now formulate the IRS that allocates resources to jobs under the constraints with the objective of minimizing the average scheduling delay. Assume we have devices $\mathbb{S}=\{s_{1},s_{2},\allowbreak...,s_{q}\}$ continuously arriving at times $\{t_{i},t_{2},\allowbreak...,t_{q}\}$. There are $m$ jobs $\mathbb{J}=\{J_{1},J_{2},\allowbreak...,J_{m}\}$ with their resource demands $\mathbb{D}=\{D_{1},D_{2},\allowbreak...,D_{m}\}$. Let $e_{ij}$ be a binary variable in the eligibility matrix, which is set to 1 if device $i$ is eligible to job $j$, and 0 otherwise. Let $x_{ij}$ be a binary variable of resource allocation, which is 1 if device $i$ is assigned to job $j$, and 0 otherwise.
We have to follow these constraints during the resource allocation:
$\sum_{j=1}^{m}x_{ij}\leq 1,\forall i\in[1,q]$
$\sum_{j=1}^{m}x_{ij}\times e_{ij}\leq 1,\forall i\in[1,q]$
$\sum_{i=0}^{q}x_{ij}=D_{j},\forall j\in[1,m]$
Therefore, the scheduling delay of each job is determined by the time it acquires the last needed device, i.e., $T_{j}=\max_{i}(x_{ij}\times t_{i})$ under these constraints. The overall objective can then be expressed as:
$\min\frac{\sum_{j=1}^{m}T_{j}}{m}$
Appendix B Theoretical Insight to the Heuristic of IRS
Lemma 1.
Given a diverse set of FL jobs with one round request, if jobs are scheduled optimally in terms of the average JCT, first within each job group and then across job groups, the resulting average JCT is optimal.
Proof.
Let us assume there is an optimal scheduling algorithm that optimizes the average JCT within each group, and there is an optimal scheduling algorithm which decides how to merge the job order across job groups to minimize the average JCT. Since the second step is assumed to provide optimal average JCT based on the previous within group job order, we only need to prove the global optimal schedule follows the order generated by the within job group step.
Venn employs smallest remaining job demand first algorithm within each job group. Since prioritizing jobs with smaller remaining resource demands has been shown to be effective in similar scheduling problems Garey et al. (1976), we skip the proof that the scheduling algorithm within job group gives the local optimal average JCT for each group.
We prove the rest by contradiction. Assume that there exists an optimal schedule $S$ that does not follow the order given by each job group. In this assumed optimal schedule $S$, let us say there are two jobs $J_{A}$ and $J_{B}$ in the same group such that $J_{A}$ comes after $J_{B}$, but $J_{A}$ has fewer resource requirements than $J_{B}$. Let us swap $J_{A}$ and $J_{B}$ to create a new schedule $S^{\prime}$. Since $J_{A}$ has fewer resource demand, the average JCT of $S^{\prime}$ will be less than that in $S$. This contradicts our original assumption that $S$ is an optimal schedule, as we’ve found a schedule $S^{\prime}$ with a lower average JCT. Therefore, the assumption is false, and the order given by each job group (sorted by resource demands) must be part of the optimal schedule. If we have an optimal scheduling across job groups, the overall average JCT will be optimal.
Appendix C Effectiveness of Venn Scheduling Heuristic
To illustrate the effectiveness of Venn’s approach, we start with proving Lemma 2, which considers a simplified case involving only two job groups with arbitrary resource contention patterns. Through mathematical proof, we can demonstrate that our algorithm achieves the optimal solution under this setting.
Lemma 2.
Given two job groups with arbitrary resource contention patterns, the scheduling plan generated by Venn as in Algorithm 1 is capable of minimizing the average scheduling delay, if a future resource allocation plan is set.
To better prove the Lemma, we introduce a new representation of the scheduling problem in a more scalable way. Firstly, as depicted in Figure 14(a), we represent the two job groups by two distinct sets of squares, where the area of each square corresponds to the size of the request demand for that job.
Secondly, to visualize the temporal dynamics of resource allocation, we refer to Figure 14(c). For the sake of this example, let’s assume a constant inflow of 100 devices per time unit. Within this set, ‘x’ devices possess memory $\geq$ 2GB, while all 100 devices have memory $\geq$ to 1GB. The y-axis is partitioned into two segments: the 0 to ‘x’ range signifies devices with memory exceeding 2GB, and the ‘x’ to 100 range represents devices with memory ranging between 1GB and 2GB.
Resource allocation over time is illustrated using rectangles, each indicating the job request to which devices are assigned. For instance, in the right subfigure of Figure 14(c), devices in the 0 to ‘x’ memory range are allocated to job group B at time 0, while those in the ‘x’ to 100 range are allocated to job group A. This representation allows us to dynamically track resource allocation across different jobs over time.
Proof.
As shown in Figure 14(a), there are $m_{A}$ requests that ask for devices with 1GB memory and $m_{B}$ jobs that request devices with 2GB memory, resulting in two job groups $A$ and $B$. The devices constantly check-in and execute one FL task, where 100% devices with memory size $\geq$ 1GB and xof the devices have memory size $\geq$ 2GB. Note that, the proof is not limited to the contention pattern draw in Figure 14(a), it can be generalized to job group with intersected resource contention and give the same conclusion.
Based on Algorithm 1, the first step is to sort these jobs within each job group by job size in ascending order (Figure 14(b)). In the second step, we generate an initial resource allocation for each job group by focusing on the job group with the scarcest resources. This results in an initial allocation plan that avoids resource sharing across job groups, setting the stage for subsequent cross-group allocations.
Based on the group-level initial allocation plan (left subfigure in Figure 14(c)), we need to determine the job order across groups, that is, to decide whether to prioritize jobs from Group $A$ over Group $B$ (right subfigure in Figure 14(c)) at current time in order to achieve a smaller average scheduling delay. In this case, we focus on determining the order of the first job with size $l$ in Group A and calculate the queuing delay difference ($\Delta t$) if we prioritize the first job from Group $A$ over Group $B$.
$\Delta t=l*m_{B}^{\prime}-(\frac{l}{1-x}-l)*m_{A}^{\prime}$
where $m_{A}^{\prime}$, $m_{B}^{\prime}$ represents the number of remaining jobs whose queuing delay may be affected by this prioritization. Since the future resource allocation is set by the previous initial allocation or assumed to be given, $m_{A}^{\prime}$, $m_{B}^{\prime}$ are feasible to get. We prioritize the first job from Group A only if $\Delta t<0$, which gives $\frac{m_{A}^{\prime}}{1-x}>\frac{m_{B}^{\prime}}{x}$, otherwise we stick with the original plan. $\Delta t<0$ is actually the prototype of the scheduling decision as in Algorithm 1 line 13.
(a) Resource supply and demand. There are 100% devices with memory size $\geq$ 1GB and xof the devices have memory size $\geq$ 2GB. There are two job groups where Group A with $m_{A}$ jobs requests for devices with memory size $\geq$ 1 GB and Group B with $m_{B}$ jobs requests for devices with memory size $\geq$ 2 GB.
(b) Step 1: Sort within Job Group.
(c) Step 2: Schedule across Job Groups.
Figure 15: Venn scheduling algorithm.
By leveraging the conclusion of Lemma 2, Venn can further generalize to the scenario with more than two job groups with arbitrary resource contention patterns. Specifically, Venn greedily compares each pair of job groups $(G_{j},G_{k})$ following the order. For each pair, Venn applies the logic proven in Lemma 2 to minimize the average scheduling delay between $G_{j}$ and $G_{k}$.
P33fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">